OpenStudy (anonymous):

3x^2=12x solve for x

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3x^2=12x }\) \(\normalsize\color{blue}{ 3x^2\color{darkgoldenrod}{-12x}=12x\color{darkgoldenrod}{-12x} }\) \(\normalsize\color{blue}{ 3x^2-12x=0 }\) \(\normalsize\color{blue}{ 3x(x^2-4)=0 }\) \(\normalsize\color{blue}{ 3x(x-2)(x+2)=0 }\)

OpenStudy (solomonzelman):

So 3x=0 or x-2=0 or x+2=0

OpenStudy (solomonzelman):

Need more help ?

OpenStudy (aum):

\[ 3x^2=12x \\ 3x^2 - 12x = 0 \\ 3x(x-4) = 0 \\ x = 0, 4 \]